### v2

# Towards Optimal Environmental Policies: Policy Learning under Arbitrary Bipartite Network Interference

Abstract
The substantial effect of air pollution on cardiovascular disease and mortality burdens is well-established. Emissions-reducing interventions on coal-fired power plants—a major source of hazardous air pollution—have proven to be an effective, but costly, strategy for reducing pollution-related health burdens. Targeting the power plants that achieve maximum health benefits while satisfying realistic cost constraints is challenging. The primary difficulty lies in quantifying the health benefits of intervening at particular plants. This is further complicated because interventions are applied on power plants, while health impacts occur in potentially distant communities, a setting known as bipartite network interference (BNI). In this paper, we introduce novel policy learning methods based on Q- and A-Learning to determine the optimal policy under arbitrary BNI. We derive asymptotic properties and demonstrate finite sample efficacy in simulations. We apply our novel methods to a comprehensive dataset of Medicare claims, power plant data, and pollution transport networks. Our goal is to determine the optimal strategy for installing power plant scrubbers to minimize ischemic heart disease (IHD) hospitalizations under various cost constraints. We find that annual IHD hospitalization rates could be reduced in a range from 20.66–44.51 per 10,000 person-years through optimal policies under different cost constraints.
Keywords: cost-constrained policy learning, environmental policy, power plants
1 Introduction
[…]
Coal-fired power plants are the largest source of sulfur dioxide emissions in the U.S., which is a major contributor to secondary PM2.5 formation (Massetti et al., 2017). A recent study found that coal-fired power plant emissions caused approximately 460,000 deaths in the U.S. Medicare population—i.e., U.S. individuals older than 65 years of age—alone from 1999-2020 (Henneman et al., 2023). An effective intervention strategy for reducing PM2.5 concentrations is installing flue gas desulfurization (FGD) equipment (so called “scrubbers”) on coal-fired power plants (Zigler et al., 2023). Scrubbers, in fact, have proven to be an effective solution, removing at least 90% of sulfur dioxide emissions from power plants on which they are applied (Srivastava et al., 2001) and, consequently, reducing the health burdens associated with exposure to power plants emissions (Zigler et al., 2023).
Given this evidence, if scrubbers were cost-free, the optimal environmental policy would be to install scrubbers at every coal-fired power plant. However, such a policy is not feasible due to the high installation and upkeep costs of scrubbers. Thus, a more nuanced intervention strategy is necessary—i.e., given a financial budget, which power plants should be ‘targeted’ for scrubbers to achieve the greatest reduction in health burdens? To illustrate this problem more clearly, consider two hypothetical power plants: power plant A is a high-emitting power plant near a densely populated, urban community with high daily temperatures (and therefore elevated susceptibility to outdoor air pollution) and power plant B is a high-emitting power plant near a sparsely populated, rural community. Based solely on this information, a scrubber in power plant A might be prioritized over power plant B. Now suppose further that pollution originating from power plant B is known to reach “downwind” densely populated urban communities due to air pollution transport patterns. Also, suppose that the cost of a scrubber at power plant B is much higher than that at power plant A. In this complex but realistic scenario, it is not clear whether power plant A or power plant B should be prioritized under a given financial budget. In the even more complex reality of hundreds of power plants on which to potentially intervene, sophisticated statistical analyses are needed to design an optimal intervention strategy.
From a statistical perspective, policy learning methods for emissions-reducing policies must accommodate two key problem features that differ significantly from more conventional causal inference settings. First, intervention strategies are implemented in coal-fired power plants (“intervention units”), but health impacts are measured in surrounding communities (“outcome units”). This creates a so-called bipartite data structure. Second, the complex process in which airborne pollutants react in the atmosphere and are transported (e.g., by the wind) means that intervening at a single power plant can potentially affect health in many distant communities, resulting in a network of connections between power plants and proximal communities. In causal inference, the presence of such connections creates interference, where one unit’s treatment can spill to (possibly many) untreated units and can affect their potential outcomes. The aggregate of these two features has been described as bipartite network interference (BNI) (Zigler & Papadogeorgou, 2018). Although a small number of papers have proposed approaches for conducting causal inference under BNI, optimal policy learning methods under BNI have not been explored to date.
1.1 Related Work
The literature on causal analysis under BNI is relatively sparse. Here we highlight primary works in the field and related works in non-BNI settings. The causal BNI literature began with Zigler & Papadogeorgou (2018), introducing estimators for causal effects under partial interference, adapting standard inverse probability weighted estimators to this simplified setting. The BNI literature expanded to accommodate more general causal estimands under experimental (Pouget-Abadie et al., 2019; Doudchenko et al., 2020) and observational settings (Zigler & Papadogeorgou, 2018; Zigler et al., 2023; Chen, Bargagli-Stoffi, Kim & Nethery, 2024), including average and heterogeneous treatment effects. More recent work considers the estimation of causal effects from longitudinal data (Chen, Bargagli-Stoffi, Kim, Henneman & Nethery, 2024; Song & Papadogeorgou, 2024). To the best of our knowledge, no work has considered the problem of policy learning in the context of BNI. The closest related works consider policy learning under interference in non-BNI settings (Zhang & Imai, 2024; Su et al., 2019; Viviano, 2024; Viviano & Rudder, 2020). Other works have considered policy learning problems under safety or budget constraints, but operate under non-BNI settings (Qiu et al., 2022; Laber et al., 2018; Xu et al., 2024). It is worth noting that with some work, one can use plug in estimates from currently established work on causal estimation (Zigler & Papadogeorgou, 2018; Doudchenko et al., 2020) in order to estimate treatment effects and optimal policy strategies under BNI. However, those methods are either not developed for observational studies, unable to rigorously accommodate arbitrary interference, and/or lack double-robustness properties for targeted policy recommendations.
1.2 Contributions and Organization
In this paper, we develop and apply methods to learn optimal power plant scrubber installation policies to minimize hospitalization burdens for ischemic heart disease (IHD) under cost constraints. Our approach will leverage a rich dataset of Medicare claims, power plant locations and features, power plant pollution transport networks, community features, and scrubber installation costs for the eastern U.S. Building methodologically on the work of Robins (2004), we propose Q-Learning and A-Learning methods for policy learning under arbitrary BNI. In contrast to much of the existing work on interference, our approach enables estimation and inference for the average treatment effects of intervening on intervention units without the need for (1) clustering assumptions or (2) causal estimands that might suffer of subjective choices of the exposure mapping such as spillover or direct effects (Hudgens & Halloran, 2008; Bargagli-Stoffi et al., 2024). We derive the asymptotic properties of these methods and demonstrate their finite sample performance via simulation studies, finding robustness of A-Learning to model misspecification in terms of estimation and inference, compared to Q-Learning. We use the proposed methods to determine the power plant scrubber installation strategy that would optimally reduce Ischemic Heart Disease (IHD) hospitalization outcomes within the United States, under cost constraints, finding that annual IHD hospitalizations could be reduced in a range from 13,353 (20.66 per 10,000 person-years) to 31,906 (44.51 per 10,000 person-years) based on different financial budget constraints.
The rest of the paper is organized as follows: Section 2 introduces our Q- and A-Learning methods and their theoretical properties. Section 3 details an extensive semi-simulated Monte Carlo study carried out to assess performance of the methods in finite samples. In Section 4, we present the details and results of our scrubber installation policy analysis. Finally, we end with discussion and limitations in Section 5. Proofs are deferred to Supp. A.
2 Methods
2.1 Setup and Objective
Setup and Notation.
Suppose we have $J$ intervention units indexed by $j\in[J]$ and $n$ outcome units indexed by $i\in[n]$. Let $\mathbf{X}^{out}_{i}$ denote the vector of covariates corresponding to outcome unit $i$, and $\mathbf{X}^{int}_{j}$ denote the vector of covariates corresponding to intervention unit $j$. $\mathbf{\mathbb{X}}_{1:J}$ will denote the covariates for all intervention units or $\mathbf{\mathbb{X}}_{1:J}=\{\mathbf{X}^{int}_{j}\}_{j\in[J]}$. $\mathbf{H}\in\mathbb{R}^{n\times J}$ or $\{\mathbf{H}_{ij}\}$ denotes the ‘interference map’ or bipartite adjacency matrix, which we will take to be known. The elements of $\mathbf{H}$ correspond to the strength of pollution transport from a particular intervention unit $j$ to each outcome unit $i$. $\mathbf{H}_{j}^{\top}=(\mathbf{H}_{1j}\dots\mathbf{H}_{ij}\dots\mathbf{H}_{nj})^{\top}$ denotes a particular column of the interference map. Similarly, $\mathbf{H}_{i}=(\mathbf{H}_{i1}\dots\mathbf{H}_{ij}\dots\mathbf{H}_{iJ})$ denotes a particular row of the interference map—that is, $\mathbf{H}_{i}$ represents the strength of connections between a particular outcome unit $i$ and each of the $J$ intervention units. $\mathbf{A}$ denotes the treatment status (or intervention status) vector. In particular, $\mathbf{A}=(A_{1}\dots A_{j}\dots A_{J})\in\{0,1\}^{J}$, where 1 corresponds to treating a particular power plant and 0 corresponds to not treating. Outcomes will depend on the exposure mapping (Aronow & Samii, 2017), which is given by $\bar{A}_{i}=\frac{1}{J}\sum_{j=1}^{J}H_{ij}A_{j}$— that is, the linear combination of treatments at the intervention unit level weighted by the strength of the connection between the outcome unit $i$ and intervention unit $j$ (interference map). Let $Y_{i}(\bar{a}_{i})$ denote the potential outcome for unit $i$ under exposure level $\bar{a}_{i}$ and $Y_{i}$ denote the observed outcome for outcome unit $i$. Note that, in this work, we will consider smaller values of $Y_{i}$ as desirable—e.g., a smaller number of IHD hospitalizations. Let the policy function for intervention unit $j$ be given by $\pi_{j}:\eta(\mathbf{X}^{out}_{i},\mathbf{H}_{j})\mapsto\Pi_{j}\subseteq[0,1]$, where $\eta$ is a functional of $\mathbf{X}^{out}_{i},\mathbf{H}_{j}$. Since we are interested in the health effects of some policy $\pi=(\pi_{1},\dots,\pi_{J})\in\Pi=[0,1]^{J}$, let $Y_{i}(\pi)$ be the potential outcome under policy $\pi$ for outcome unit $i$, similar to how ‘$Y(\rho)$’ is defined in Section 2 of Qiu et al. (2022). Finally, let $e_{j}$ be the propensity score for intervention unit $j$, or
$e_{j}=\mathbb{P}(A_{j}=1\mid\mathbf{X}^{int}_{j}).$
Objective.
The objective is to determine the optimal treatment policy $\pi^{*}\in\Pi=[0,1]^{J}$ which minimizes $\mathbb{E}[Y_{i}(\pi)]$, the mean outcome under policy $\pi$— e.g. the treatment policy that yields the smallest number of IHD hospitalizations.
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi}\mathbb{E}[Y_{i}(\pi)]$ (1)
2.2 Nonparametric Identification
We state the identification assumptions required to identify our estimand Equation 1 from data.
Identification Assumptions
  1. (A1)
    Consistency of outcome units: $Y_{i}=Y_{i}(\bar{a}_{i})=Y_{i}(\mathbf{A})$.
  2. (A2)
    Positivity: $P(\bar{a}_{i}\mid\mathbf{X}^{out}_{i})>0$.
  3. (A3)
    Unconfoundedness and Intervention Unit Unconfoundedness: $Y_{i}(\bar{a}_{i})\perp\bar{a}_{i}\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J}$.
  4. (A4)
    Intervention covariates are independent of potential outcomes given the exposure mapping $\bar{a}_{i}$ and outcome covariates: $Y_{i}(\bar{a}_{i})\perp\mathbf{\mathbb{X}}_{1:J}\mid\mathbf{X}^{out}_{i},\bar{a}_{i}$.
Assumptions (A1)-(A3) are adaptations of the standard consistency, positivity, and ‘no unmeasured confounding’ assumptions in causal inference to our BNI setting. Note that the second equality of (A1) is made since $\mathbf{H}$ cannot be controlled and we only have control over $\mathbf{A}$. Assumption (A4) encodes domain knowledge from air pollution policy: outcome unit potential outcomes are determined by aggregate exposure levels $\bar{A}_{i}$ rather than individual treatment values of intervention units $A_{j}$. Thus, when learning the distribution of $Y_{i}(\bar{a}_{i})$ given $\bar{A}_{i}$ and $\mathbf{X}^{out}_{i}$, it is not necessary to condition on $\mathbf{\mathbb{X}}_{1:J}$. While this assumption is not strictly necessary, it enables simplifications of the proposed modeling approaches by leveraging domain knowledge.
Proposition 2.1 (Identification).
Under (A1)-(A4), we have
$\mathbb{E}[Y_{i}(\bar{a}_{i})]=\mathbb{E}\big{[}\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]\big{]}$ (2)
This means that we can instead consider the following objective function
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi}\mathbb{E}_{n}\big{[}\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]\big{]}$ (3)
Note that, in the real world, we may have cost constraints—e.g., a total budget for installation cost of scrubbers. Let $C$ denote the total cost constraint and for $j\in[J],$ let $c_{j}=\mbox{Cost}(A_{j}=1)-\mbox{Cost}(A_{j}=0)$ denote the cost difference between treating and non treating the intervention unit $j$. To enable identification of cost-constrained optimal policies, we require these costs $c_{j}$ to be known.
Cost Assumption
  1. (A5)
    Known treatment costs: For $j\in[J]$, $c_{j}$ is known and $c_{j}\geq 0$.
Corollary 1 (Cost-constrained Optimal Policy).
Assume (A1)-(A4) and (A5). Then, the optimal cost constrained optimization problem is given by:
[…]
2.3 Optimal Policy Learning using Parametric, Linear Exposure Models
In this section, we propose Q-Learning and A-Learning methods for solving Equation 4. Motivated by domain knowledge that outcome unit potential outcomes are determined by an aggregate exposure level, we assume the outcome model is linear in $\bar{A}_{i}$. For illustration purposes, we present methods based on parametric forms. Please note that such results can be generalized. Namely, the methods below can be extended to multiple timepoints, semi/nonparametric models under appropriate smoothness conditions, and general outcome model forms. See Supp. F for details. Here, we assume the following outcome model.
Modeling Assumption
  1. (A6)
    Outcome model is linear in the exposure mapping:
    $Y_{i}(\bar{a}_{i};\boldsymbol{\theta}_{0})$ $=\mu(\bar{a}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\theta}_{0})+\epsilon_{i}$
    $=f_{0}(\mathbf{X}^{out}_{i},\boldsymbol{\alpha}_{0})+\bar{a}_{i}\cdot f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{0})+\epsilon_{i}$ (5)
for $\boldsymbol{\theta}_{0}=(\boldsymbol{\alpha}_{0},\boldsymbol{\beta}_{0})\in\Theta\subseteq\mathbb{R}^{|\boldsymbol{\theta}_{0}|}$, which parameterize $f_{0}(\mathbf{X}^{out}_{i};\boldsymbol{\alpha}_{0})$ and $f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta}_{0})$ respectively, and mean zero independent random variable $\epsilon_{i}$. Above, $f_{0}$ captures the main effect of the covariates and $f_{A}$ captures the treatment effect. By Equation 5, we can express the objective function as follows:
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i};\boldsymbol{\theta}_{0}]$ $=\frac{1}{n}\sum_{i=1}^{n}f_{0}(\mathbf{X}^{out}_{i};\boldsymbol{\alpha}_{0})+\frac{1}{n}\sum_{i=1}^{n}\bar{A}_{i}f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta}_{0})$
$=\frac{1}{n}\sum_{i=1}^{n}f_{0}(\mathbf{X}^{out}_{i};\boldsymbol{\alpha}_{0})+\frac{1}{n}\sum_{j=1}^{J}A_{j}\frac{1}{J}\sum_{i=1}^{n}H_{ij}f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta}_{0})$
Let
$\texttt{TotalEffect}_{j}=\frac{1}{J}\sum_{i=1}^{n}H_{ij}f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta}_{0}),\quad j\in[J]$ (6)
Hence the effect of treating an intervention unit $j$ is fully captured by $\texttt{TotalEffect}_{j}$, precluding the need for defining a direct or spillover effect or making clustered interference assumptions. The quantity in Equation 6 will serve as a key object of interest throughout our analyses below.
Optimal Policy under a Linear Exposure Model.
Since smaller values of $Y$ are preferred (e.g. less hospitalizations), the optimal regime $\pi^{*}$ in the absence of cost constraints would be given by
$\pi_{j}^{*}(\texttt{TotalEffect}_{j})=\textbf{1}\{\texttt{TotalEffect}_{j}<0\}$ (7)
We are now set up to state our Q-Learning and A-Learning approach to obtaining statistical inference procedures. At a high-level, Q-Learning and A-Learning posit regression models for the mean outcome and determine the optimal policy by using a plug-in estimate for the decision rule. Q-Learning relies on a correctly-specified outcome regression model, while A-Learning is a variant of Q-Learning that fits a modified estimating equation to endow double robustness (Schulte et al., 2014). In our case of Equation 5, Q-Learning requires $\mu=(f_{0},f_{A})$ to be correctly specified to obtain the optimal policy and inference. As we will see in our approach below, A-Learning requires $f_{A}$ to be correctly specified but only requires correct specification of only one of baseline model $f_{0}$ or propensity score model $e$ to obtain valid inference and the optimal policy.
2.3.1 Q-Learning
Parametric Q-Learning proceeds by proposing a model $\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i};\boldsymbol{\theta}_{0}]$, and solving Equation 4 using the plug-in estimate $\hat{\boldsymbol{\theta}}$.
Estimation.
To estimate $\boldsymbol{\theta}_{0}$, we perform least squares regression, i.e. we solve the following estimating equations:
$\frac{1}{n}\sum_{i=1}^{n}\phi(Y_{i};\boldsymbol{\theta})=0$ (8)
where $\phi(Y_{i};\boldsymbol{\theta})=(Y_{i}-\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\theta}))d_{i}(\boldsymbol{\theta})$ and $d_{i}(\boldsymbol{\theta})=\frac{\partial\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}$.
Inference.
Under suitable regularity conditions, we find for covariance matrix $\Sigma$ defined in Section A.2,
[…]
2.3.2 A-Learning
Q-Learning requires the full model $\mu=(f_{0},f_{A})$ to be correctly specified in order to obtain the optimal policy and inference. In this section, we propose a doubly robust estimator for $\boldsymbol{\beta}$. Below, we are explicit about the dependence on $(\boldsymbol{\alpha},\boldsymbol{\beta})$ since they are estimated using different estimating equations. Let the propensity score for intervention unit $j$, $e_{j}$, be parameterized by $\boldsymbol{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{|\boldsymbol{\gamma}_{0}|}$.
$e_{j}=\mathbb{P}(A_{j}=1\mid\mathbf{X}^{int}_{j};\boldsymbol{\gamma}_{0}).$ (9)
Estimation
We propose to estimate $\boldsymbol{\beta}_{0}$ by solving for the roots of the following estimating equation:
$\Phi_{n}(\boldsymbol{\beta},\boldsymbol{\alpha},\boldsymbol{\gamma})$ $=\frac{1}{n}\sum_{i=1}^{n}\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})(Y_{i}-\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\alpha},\boldsymbol{\beta}))(\bar{A}_{i}-\widehat{\bar{A}}_{i})$ (10)
where $\widehat{\bar{A}}_{i}=\frac{1}{J}\sum_{j=1}^{J}e_{j}(\mathbf{X}^{int}_{j},\hat{\gamma})H_{ij}$ and $\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})$, following the notation of Eq. 31 in (Schulte et al., 2014), is an arbitrary function that is of dimension $|\boldsymbol{\beta}|$.
We solve this estimating equation jointly with the following estimating equations for $\bm{\alpha_{0}},\bm{\gamma_{0}}$. For $\alpha$, we solve
$\frac{1}{n}\sum_{i=1}^{n}\frac{\partial}{\partial\boldsymbol{\alpha}}f_{0}(\mathbf{X}^{out}_{i},\boldsymbol{\alpha})(Y_{i}-\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\alpha},\boldsymbol{\beta}))=0$ (11)
For $\gamma$, we solve
$\frac{1}{J}\sum_{j=1}^{J}\frac{\partial}{\partial\boldsymbol{\gamma}}e(\mathbf{X}^{int}_{j},\boldsymbol{\gamma})(A_{j}-e(\mathbf{X}^{int}_{j},\boldsymbol{\gamma}))=0$ (12)
Remark 2.1.
In Equation 10, the choice of $\lambda$ is flexible. For example, one can take this to be $\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})=c_{i}\frac{\partial}{\partial\boldsymbol{\beta}}f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta})$ for $c_{i}\in\mathbb{R}$. This is similar to (Leon et al., 2003; Robins, 2004; Schulte et al., 2014), in causal inference literature on defining a ‘class’ of estimating equations that may satisfy consistent estimation, asymptotic normality, and double-robustness. During our simulations and analysis, we take $c_{i}=\frac{1}{J}\sum_{j=1}^{J}H_{ij}$.
This formulation yields an approach to estimating $\boldsymbol{\beta}$ that is doubly-robust, as shown by the lemma below; if either $e$ or $f_{0}$ is correctly specified, we have consistent parameter estimation for $\beta$ (Bang & Robins, 2005). We clarify that the following approach still requires $f_{A}$ to be correctly specified.
Lemma 1 (Double Robustness).
If $e$ or $f_{0}$ is correctly specified, then under the identification conditions (A1)-(A4),
$\mathbb{E}[\Phi_{n}(\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\gamma}})]=0$
Inference
Theorem 1 (Asymptotic Normality).
Under suitable regularity conditions and (A11), that $\frac{J}{n}\rightarrow\mathcal{R}$ a fixed constant, as $J,n\rightarrow\infty$, we have for covariance matrix $\Omega$ defined in Section A.3.1,
[…]
2.4 Estimating the Optimal Policy
With the treatment effect estimates in hand via $\hat{\boldsymbol{\beta}}$ from either Q-Learning or A-Learning, we now describe how to estimate the optimal policy.
Unconstrained Optimal Policy.
From Equation 7, we use the plug-in $\hat{\boldsymbol{\beta}}$ to find:
$\hat{\pi}_{j}^{*}(\widehat{\texttt{TotalEffect}}(\hat{\boldsymbol{\beta}}))=\textbf{1}\{\widehat{\texttt{TotalEffect}}(\hat{\boldsymbol{\beta}})<0\}$ (13)
Cost-Constrained Optimal Policy.
The constrained optimal policy is given by solving Equation 4 with plug-in $\hat{\boldsymbol{\beta}}$. If we assume that the costs and cost constraints are positive (i.e. $c_{j},C>0$), we can obtain an efficient $O(J)$ computable solution to Equation 4 via the fractional knapsack solution, taking $c_{j}$ as weights and $-Y_{i}$ as profits (see Ch. 15 Cormen et al. (2009)). To translate this to our setup, we define the following:
$\texttt{BenefitCost}_{j}=\frac{\texttt{TotalEffect}_{j}}{c_{j}}$ (14)
This quantity represents the benefit-cost ratio of installing a scrubber at a power plant $j$ relative to the cost. Such “relative benefits” play a key role in determining cost-constrained optimal intervention strategies, as $\texttt{TotalEffect}_{j}$ does for the unconstrained counterpart. We will estimate this ratio using our model output to explain and contextualize our findings in Section 4.4.2.
3 Simulation Study
In this section, we conduct an empirical Monte Carlo study to validate the methods in Section 2. The covariates and interference matrix are the same as from our data application across all simulations. Next to these “empirical variables”, we randomly generate treatment assignments and outcomes according to a fixed Signal-to-Noise Ratio (SNR). The Bias of $\boldsymbol{\beta}$ (Equation C.1), the RMSE of $\widehat{\texttt{TotalEffect}}$ (Equation C.2), and the coverage probabilities are reported across all simulations. The equations for the metrics and parameters used for simulation are found in Supp. C. Throughout simulations, we assume the following form for the true outcome and propensity score models:
Outcome Model.
We assume that $f_{0},f_{A}$ are quadratic in $\mathbf{X}^{out}_{i}$, where $\alpha_{\mathrm{intercept}}\in\mathbb{R}$ and $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{\mathrm{dim}(\mathbf{X}^{out}_{i})}$. Similarly, $\beta_{\mathrm{intercept}}\in\mathbb{R}$ and $\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2}\in\mathbb{R}^{\mathrm{dim}(\mathbf{X}^{out}_{i})}$.
$f_{0}(\mathbf{X}^{out}_{i},\alpha)=\alpha_{\mathrm{intercept}}+\mathbf{X}^{out}_{i}\boldsymbol{\alpha}_{1}+(\mathbf{X}^{out}_{i})^{2}\boldsymbol{\alpha}_{2}$ (15)
$f_{A}(\mathbf{X}^{out}_{i},\beta)=\beta_{\mathrm{intercept}}+\mathbf{X}^{out}_{i}\boldsymbol{\beta}_{1}+(\mathbf{X}^{out}_{i})^{2}\boldsymbol{\beta}_{2}$ (16)
Propensity Score Model.
We model $e_{j}$ using a logistic regression model where $\gamma_{\mathrm{intercept}}\in\mathbb{R}$ and $\boldsymbol{\gamma}_{1},\boldsymbol{\gamma}_{2}\in\mathbb{R}^{\mathrm{dim}(\mathbf{X}^{out}_{i})}$.
$\mathrm{log}\left(\frac{e_{j}}{1-e_{j}}\right)$ $=\gamma_{\mathrm{intercept}}+\mathbf{X}^{int}_{j}\boldsymbol{\gamma}_{1}+(\mathbf{X}^{int}_{j})^{2}\boldsymbol{\gamma}_{2}$ (17)
In the simulations, we consider scenarios with fully correct model specification and with $f_{0}$ and/or $e$ misspecified to demonstrate the double robustness property of A-learning. Misspecification of each component is induced by fitting models imposing linear relationships with covariates in a given component instead of quadratic. Specifically, outcome model misspecification is carried out by setting $\boldsymbol{\alpha}_{2}=0$. The misspecification of the propensity score model is introduced by setting $\boldsymbol{\gamma}_{2}=0$. Throughout, to construct an empirically grounded simulation framework that accurately reflects the real-world dynamics of our data application, $\boldsymbol{\gamma}$ is chosen so that simulated average of propensities $\frac{1}{J}\sum_{j=1}^{J}e_{j}$ is within 0.01 of the observed empirical average of treatments in the dataset (see final row of Table B.1). Further, $\boldsymbol{\theta}=(\boldsymbol{\alpha},\boldsymbol{\beta})$ is chosen so that the simulated average $\frac{1}{n}\sum_{i=1}^{n}Y_{i}$, based on the expected exposure level $\bar{A}=\frac{1}{J}\sum_{j=1}^{J}H_{ij}e_{j}$ from Step 2, is within 0.001 of the empirical average in the dataset (see penultimate row of Table B.2).
Our simulation setup is as follows.
  1. 1.
    Preprocess data: Standardize the covariates of $\mathbf{X}^{out}_{i},\mathbf{X}^{int}_{j}$.
  2. 2.
    Generate intervention unit treatments: $A_{j}\sim\mathrm{Bernoulli}(e_{j})$, where $e_{j}$ follows Equation 17.
  3. 3.
    Compute exposure mapping: $\bar{A}_{i}=\frac{1}{J}\sum_{j\in[J]}H_{ij}A_{j}$
  4. 4.
    Generate outcomes: $Y_{i}(\bar{A}_{i})$ such that the SNR is 3, i.e.
    $Y_{i}=f_{0}(\mathbf{X}^{out}_{i},\boldsymbol{\alpha})+\bar{A}_{i}\cdot f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta})+\epsilon_{i}$
    where $\mathbb{V}(\epsilon_{i})=\mathbb{V}\left(\frac{\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i};\boldsymbol{\theta}]}{3}\right)$ and $f_{0}$, $f_{A}$ follow Equation 15 and 16 respectively.
  5. 5.
    Estimate parameters: $\boldsymbol{\theta}=(\boldsymbol{\alpha},\boldsymbol{\beta})$ using each of the following:
    • Q-learning with correctly specified outcome model $\mu=(f_{0},f_{A})$
    • Q-learning with misspecified $f_{0}$, correctly specified $f_{A}$
    • A-learning with correctly specified outcome $\mu=(f_{0},f_{A})$ and propensity score $e$
    • A-learning with correctly specified outcome $\mu=(f_{0},f_{A})$ and misspecified $e$
    • A-learning with $f_{0}$ misspecified but $f_{A}$ and $e$ correctly specified
    • A-learning with both $f_{0}$ and $e$ misspecified but $f_{A}$ correctly specified
    where the Q-learning models are estimated using least squares regression and the A-learning models solve Equation 11, 12 and 10 jointly.
  6. 6.
    Iterate: Repeat steps (2-5) 1000 times.
    1. (i)
      Compute Bias, RMSE, and the average empirical 95% confidence interval coverage probabilities for $\hat{\boldsymbol{\beta}}$ (across all the dimensions of $\mbox{dim}(\mathbf{X}^{out}_{i})$).
Simulation Results
Method BS ($f_{0}$) PS ($e$) Bias RMSE Coverage
Q-Learning - 0.021 0.73 94.84
Q-Learning - 0.22 7.10 15.40
A-Learning 0.024 1.32 98.57
A-Learning 0.024 1.32 98.37
A-Learning 0.057 1.74 95.23
A-Learning 0.058 1.75 94.56
Table 1: Simulation Results over 1000 iterations. BS denotes baseline model, $f_{0}$, and PS denotes propensity score model, $\pi$. See the definitions for Bias (Equation C.1) and RMSE (Equation C.2) above.
The results from 1000 simulations are shown in Table 1. The correctly specified Q-Learning model achieves the lowest bias and RMSE, consistent with common knowledge that a correctly-specified outcome model tends to outperform other methods. A-Learning with correctly specified models is competitive with Q-Learning correctly specified. Under misspecification of the propensity score model ($e$), A-Learning gives nearly equivalent performance to A-Learning under a fully correct model specification (first and second rows of A-Learning). As we introduce the misspecification of the outcome model in Q-Learning and A-Learning, the error metrics increase.
Based on these simulations, under equivalent misspecification forms, the methodology seems less sensitive to propensity score model misspecification than outcome model misspecification. This phenomenon was also observed in simulations on doubly robust causal effect estimation under BNI in Chen, Bargagli-Stoffi, Kim & Nethery (2024). Regardless, under equivalent baseline model misspecification (second row of Q-Learning and third row of A-Learning), A-Learning results in lower estimation error than Q-Learning. A-Learning yields good empirical coverage for significance level $\alpha=0.05$, even with misspecified models. In contrast, under equivalent outcome model misspecification, Q-Learning has extremely low coverage, affecting inference more severely than A-Learning.
In general, we find a lower estimation error in treatment effect estimates and a higher coverage under various levels of misspecification in A-Learning compared to misspecified Q-Learning, validating the theoretical results shown above in Section 2.3.1 and Section 2.3.2.
4 Optimal Power Plant Scrubber Installation Policy
We applied the optimal policy methods developed above to real data to determine the allocation of power plant scrubbers that minimizes the hospitalization rates for IHD among Medicare beneficiaries in the United States given a financial budget. The data, analysis of policy learning, and the results are described below. Additional analyses on non-linear outcome model fitting and equivalent count analysis are found in the Supplement.
4.1 Data Description
Our data consist of three parts: (i) outcome level data with Medicare beneficiary information, (ii) intervention level data with power plant characteristics and scrubber cost information, and (iii) the interference map or characterization of air pollution transport effects between outcome units and intervention units via a climate model known as HYSPLIT Average Dispersion (HyADS) (Henneman et al., 2019). We detail each component below.
4.1.1 Outcome Level Data: Medicare Beneficiary Data
The outcome data are derived from Medicare inpatient claims data for all 27,312,190 fee-for-service beneficiaries residing in the U.S. in 2005. The health outcome of interest is the ZIP code-level ischemic heart disease (IHD) inpatient hospitalization rate per 10,000 person-years in 2005 among these Medicare beneficiaries. For each ZIP code, U.S. Census socioeconomic and demographic features (from the 2000 decennial census), meteorological (Kalnay et al., 1996), and smoking rate (Dwyer-Lindgren et al., 2014) covariate data were obtained. Further, due to the relatively small number of power plants in the Western U.S., we remove ZIP codes in the U.S. states of Wyoming, Colorado, Idaho, Utah, Arizona, New Mexico, Nevada, California, Oregon, Montana, and Washington. After initial cleaning, we have $n=26,674$ ZIP codes. The ZIP code covariates used in our analysis, and corresponding descriptive statistics, are given in Table B.2.
The ZIP code-level IHD hospitalization rate per 10,000 person-years is used as the outcome in our statistical models, in order to account for different amounts of person-time of exposure across ZIP codes. Our primary results shown below are based on minimizing rates of hospitalization. The results stated in terms of total counts of hospitalizations are detailed in Section E.3.
4.1.2 Intervention Level Data: Power Plant Data
Data on scrubber status in the year 2005 and plant characteristics for $J=409$ coal–fired power plants concentrated in the Eastern US, which serve as our intervention units of interest, were obtained from the U.S. Environmental Protection Agency Air Markets Program Database. The plant covariates used in our analyses, along with descriptive statistics, are given in Table B.1. Scrubber installation cost information for power plants that had a scrubber installed by 2021 was obtained from the Energy Information Administration (EIA) website.
4.1.3 Interference Map: HyADS Climate Model
To characterize the interference structure or the bipartite adjacency matrix, we utilize the HyADS model. HyADS is a pollution transport model that yields a unit-less metric quantifying the amount of emissions from an individual power plant that were transported (e.g., by wind) to a particular ZIP code (Henneman et al., 2019). These values, which were calculated for all power plant pairs and ZIP codes in the data, form the elements of $H$.
4.2 Data Preprocessing
We perform log transformations of several covariates (TotHeatInput, TotPop, and TotOpTime) in order to reduce skew in their distribution and standardization of covariates from non-categorical outcome and intervention units. Prior to final analysis for A-Learning, to reduce instability we perform trimming of intervention units with low estimated probability of treatment (i.e., very small estimated propensity scores) (Stürmer et al., 2010). Specifically, we trim below the bottom 5th percentile of estimated propensity scores (approximately $0.018$). After trimming intervention units, we end up with $J=388$ power plants.
4.3 Cost Model Development
In order to find the optimal allocation of scrubbers under cost-constraints, we need to quantify the cost of installing a scrubber at each power plant. Because only a subset of plants has scrubbers installed (and therefore has observed installation cost information), we build a model to impute scrubber installation costs for power plants without scrubbers based on covariates available from our intervention dataset. Since the cost data from the EIA include information from various stacks within a given power plant facility, we define the cost ($c_{j}$ in Equation 4) to be the average cost of installing a scrubber at a given power plant.
The EIA data report scrubber installation costs for 135 of the power plants represented in our data. These 135 samples serves as our training set. We follow a standard machine learning approach to build our cost model. First, we split our training data ($J=135$) into a (sub) training and validation set using an 80/20 split, resulting in ($J=108$) and ($J=27$) respectively. Then, we test three models—namely, random forest, support vector regression, and linear regression—on the validation set. The model with the best performance is selected. Finally, the selected model is then trained on our full training data, and used to generate scrubber installation cost predictions on power plant data with missing cost information ($J=225$). Performance metrics on the validation set, variables’ importance metrics from our final random forest model, and cost predictions are shown in Section D.2. Under our cost model, the total cost for installing a scrubber at all of the 388 power plants is approximately \$43,969,292,000.
4.4 Policy Analysis
We apply the approaches developed above to (1) quantify the effects of FGD scrubber installation at each power plant on IHD hospitalizations in 2005 and (2) identify optimal scrubber installation policies under cost constraints. For our primary analysis, we use A-Learning to protect against model misspecification.
4.4.1 TotalEffect Rates Analysis
For interpretability, we build an outcome model and a propensity score model akin to Equation 15 and Equation 16, respectively, but linear in covariates instead of quadratic. Sensitivity analyses with more complex models are reported in Section E.2 and support the conclusions of the linear model with larger IHD reductions found. Upon fitting the A-Learning model, we find that $100\%$ of power plants have estimated protective effects if treated, indicating that scrubber installation would reduce IHD hospitalizations. Graphically Fig. 1 shows a plot of $\widehat{\texttt{TotalEffect}}_{j}$ on the U.S. map and in a histogram. Installing scrubbers at most power plants tend to have small effects, save for a handful of influential power plants in the Midwest. For example, consider the power plant that has the greatest estimated reduction in IHD hospitalization rates if a scrubber in installed, found in Tennessee. If we install a scrubber at this power plant, we would find a 0.61 (408) reduction in IHD hospitalization rate per 10,000 person-years (counts) in 2005 among Medicare beneficiaries. An exploratory analysis testing $H_{0}:\texttt{TotalEffect}_{j}\geq 0,H_{a}:\texttt{TotalEffect}_{j}<0$ based on Theorem 1 yields evidence of statistically significant protective population effects when installing scrubbers at the majority of the power plants for significance level $\alpha=0.05$; Fig. E.1 shows $\sim 95\%$ of p-values are statistically significant. See Supplementary Materials for more information, including p-values and 95% CIs.
Refer to caption Refer to caption
(a) (b)
Figure 1: (a) displays the plot of $\widehat{\texttt{TotalEffect}}_{j}$ on a U.S. map. (b) displays the histogram of $\widehat{\texttt{TotalEffect}}_{j}$.
4.4.2 Optimal scrubber policies under cost constraints
Because, to our knowledge, policymaking entities have not specified exact budgets for scrubber installations, here we consider various total cost budgets defined as percents (10%, 20%,…,90%) of the estimated cost of installing scrubbers at all power plants used in our analysis (\$43,969,292,000, as reported above). Under each cost budget, the optimal treatment policy, i.e., the set of plants that are selected for scrubber installation, and the subsequent average reduction in IHD hospitalizations per 10,000 person years are shown in Fig. 2. Equivalent analysis for counts are shown in Section E.3. Under a budget constituting 10% of the cost of universal scrubber installations at all US power plants, annual IHD hospitalizations can be reduced by 13,353 (20.66 per 10,000 person-years) via the optimal scrubber installation policy. The plants selected for scrubbers under this budget are largely concentrated in the Ohio River Valley and Midwest. Under a budget constituting 90% of universal installation costs, 31,447 (44.25 per 10,000 person-years) IHD hospitalizations can be avoided annually. As shown in Fig. 2, the health benefits of additional scrubber installations seems to level off for budgets above approximately 60% of universal installation cost, and the biggest reductions occur as the budget increases from 10% – 30% of universal cost.
To investigate this, we map the $\widehat{\texttt{TotalEffect}}_{j}$ alongside the $\widehat{\texttt{BenefitCost}}_{j}$ in Fig. 3. The scale of $\widehat{\texttt{BenefitCost}}_{j}$ is much smaller, so we negate and log transform it for visualization. This enables us to compare how ‘dark’ different power plants are in Fig. 3, to give more insights into the factors influencing selection for treatment in the optimal policies. As we can see, the power plants around the Midwest that have the greatest reductions in $\widehat{\texttt{TotalEffect}}_{j}$ generally have competitive cost-benefit ratios. Interestingly, we can observe that several other power plants in the surrounding regions are also prioritized in the cost-constrained intervention strategies because, although their $\widehat{\texttt{TotalEffect}}_{j}$ is relatively small, they also have cheaper scrubber installation costs, leading to a more competitive $\widehat{\texttt{BenefitCost}}_{j}$.
We end with a comparison of the cost-constrained policies from Fig. 2 to a more intuitive policy that policymakers may carry out under realistic cost constraints. Recall from Section 2.4 that the optimal policy to Equation 4 is obtained from the fractional knapsack solution. Informally, this means that one must: (a) estimate $\texttt{BenefitCost}_{j}$, and (b) treat power plants $j$ with the lowest $\widehat{\texttt{BenefitCost}}_{j}$ (as lower hospitalizations are better) so long as the cost constraint is satisfied. Intuitively, a policymaker may do the same as above, but instead treat power plants with the lowest $\widehat{\texttt{TotalEffect}}_{j}$ (i.e. the greatest reduction in IHD hospitalizations) rather than the lowest $\widehat{\texttt{BenefitCost}}_{j}$; Fig. 4 shows the effect of enacting such a policy. Comparing Fig. 2 to Fig. 4, we find uniformly lower reductions (i.e. greater health benefits) from solving Equation 4 compared to a more intuitive solution.
Refer to caption Refer to caption Refer to caption
(a) 10% Total Cost (b) 20% Total Cost (c) 30% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 40% Total Cost (b) 50% Total Cost (c) 60% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 70% Total Cost (b) 80% Total Cost (c) 90% Total Cost
Figure 2: Grid with the reduction of IHD Hospitalizations/10,000 Person Years, varying the spending from 10%-90% of budget
Refer to caption Refer to caption
(a) (b)
Figure 3: (a) displays the plot of $-\widehat{\texttt{TotalEffect}}_{j}$ on a U.S. map, colored by intensity of $\widehat{\texttt{TotalEffect}}_{j}$. (b) displays an analagous plot for $-\log(-\widehat{\texttt{BenefitCost}}_{j})$.
Refer to caption Refer to caption Refer to caption
(a) 10% Total Cost (b) 20% Total Cost (c) 30% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 40% Total Cost (b) 50% Total Cost (c) 60% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 70% Total Cost (b) 80% Total Cost (c) 90% Total Cost
Figure 4: Grid with the reduction of IHD Hospitalizations/10,000 Person Years, varying the spending from 10%-90% of budget, optimizing based on $\widehat{\texttt{TotalEffect}}$ rather than $\widehat{\texttt{BenefitCost}}$.
5 Discussion
Understanding how to optimally assign an intervention is critically important to optimize policy effects while keeping the costs fixed. In this paper, we proposed a new methodology to produce optimal policies in the context of bipartite network interference. Using our proposed methodology, we carry out the first statistical policy learning analyses for air pollution policy, finding up to 31,906 (44.51 per 10,000 person-years) IHD hospitalizations can be avoided annually among Medicare fee-for-service beneficiaries. This analysis supports previous findings that reducing pollutants from large-scale emissions reduces public health burden (Henneman et al., 2023). The analysis of $\widehat{\texttt{BenefitCost}}$ sheds light onto how nuanced this analysis can be under realistic cost constraints. Different budgets and policy choices can yield variable effects on public health impact.
From a methodological standpoint, we proposed the first BNI policy learning methodology based on Q-Learning and A-Learning under arbitrary BNI and validated our proposed methods in an extensive simulation study that closely mimics our real world setting. Our approach does not require clustering assumptions or defining direct and spillover estimands as commonly done in interference literature; such assumptions are untenable in this problem due to air pollution transport. The proposed methodologies hold general applicability for future large environmental policy studies, as it can encompass variants of clustered BNI.
There are several limitations to note. We conduct our primary analysis assuming a linear exposure mapping and that HyADS is a primary effect modifier for power plant interventions. Such an assumption enables computationally feasible modeling and aligns with domain knowledge, but it is untestable in practice. Methodological limitations of this work are assuming costs $c_{1}\dots c_{J}$ are known and the interference map $\mathbf{H}$ is known (as in practice HyADS is an estimation method). Regardless, assuming that $\mathbf{H}$ is known, is a common assumption made in bipartite literature (see, e.g., Papadogeorgou et al., 2019; Doudchenko et al., 2020; Chen, Bargagli-Stoffi, Kim & Nethery, 2024; Chen, Bargagli-Stoffi, Kim, Henneman & Nethery, 2024). Future work should address the limitations above. Further, given recent evidence of disproportionate burden to marginalized subgroups in air pollution regulatory policy (Zeger et al., 2008; Hajat et al., 2015; Kioumourtzoglou et al., 2015; Di et al., 2017; Jbaily et al., 2022; Josey et al., 2023; Chen, Bargagli-Stoffi, Kim & Nethery, 2024), extending this work to ensure ‘fair’ policy learning (Viviano & Bradic, 2024) stands as an important future direction.
Funding.
This work was funded by the Health Effects Institute award 4998-RFA22-2/24 and NIH grants T32ES714242, K01ES032458.
SUPPLEMENTARY MATERIAL TO
“Towards Optimal Environmental Policies: Policy Learning under Arbitrary Bipartite Network Interference”
Raphael C. Kim1, Falco J. Bargagli-Stoffi2, Kevin L. Chen1, Rachel C. Nethery1
1Department of Biostatistics, Harvard T.H. Chan School of Public Health (Boston, MA)
2Department of Biostatistics, UCLA Fielding School of Public Health (Los Angeles, CA)
Appendix A Proofs
The results below employ the same strategy used in Su et al. (2019). However, we require results in a setting where there is arbitrary bipartite network interference. Since the setup is unique, we explicitly write out the results.
A.1 Proof of Proposition 2.1
$\mathbb{E}[Y_{i}(\bar{a}_{i})]$ $=\mathbb{E}[\mathbb{E}[Y_{i}(\bar{a}_{i})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J}]]$
=(A3)${{=}}\mathbb{E}[\mathbb{E}[Y_{i}(\bar{a}_{i})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\bar{A}_{i}]]$
=(A4)${{=}}\mathbb{E}[\mathbb{E}[Y_{i}(\bar{a}_{i})\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]]$
=(A1)${{=}}\mathbb{E}[\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]]$
A.2 Proof of Q-Learning Inference
Similar to the standard assumptions for M-estimation (Stefanski & Boos, 2002), the following are required for statistical inference of $\mathbb{E}[Y_{i}(\bar{a}_{i};\boldsymbol{\theta}_{0})]$ under Q-Learning:
Inference Assumptions
  1. (A7)
    The outcome covariates, intervention covariates, and interference map are bounded. $||\mathbf{X}^{out}_{i}||_{\infty}<\infty$, $||\mathbf{X}^{int}_{j}||_{\infty}<\infty,||\mathbf{H}||<\infty$
  2. (A8)
    Parameters $\boldsymbol{\theta}$ lie in a compact space $\mathbf{\boldsymbol{\theta}}$ with $\boldsymbol{\theta}_{0}$ as an inner point.
  3. (A9)
    $\Sigma_{d}=\lim\limits_{n\rightarrow\infty}\frac{1}{n}\sum_{i=1}^{n}[-\frac{\partial\phi(Y_{i};\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}}]$ is finite and non degenerate
  4. (A10)
    $\Sigma_{\phi}=\lim\limits_{n\rightarrow\infty}\frac{1}{n}\sum_{i=1}^{n}\phi(Y_{i};\boldsymbol{\theta}_{0})\phi(Y_{i};\boldsymbol{\theta}_{0})^{\top}$ is positive semi-definite.
Proof.
Using Taylor’s theorem as in Section 2 of (Stefanski & Boos, 2002):
$\sqrt{n}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0})$ $=[-\frac{1}{n}\sum_{i=1}^{n}\frac{\partial\phi(Y_{i};\boldsymbol{\theta})}{\partial\boldsymbol{\theta}^{\top}}|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}}]^{-1}[\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi(Y_{i};\boldsymbol{\theta}_{0})]+o_{p}(1)$
[…]
Consistent variance estimation follows by $\hat{\Sigma}=\Sigma(\hat{\boldsymbol{\theta}})=\frac{1}{n}\Sigma_{d}^{-1}(\hat{\boldsymbol{\theta}})\Sigma_{\phi}(\hat{\boldsymbol{\theta}})(\Sigma_{d}^{-1}(\hat{\boldsymbol{\theta}}))^{\top}$ for $\Sigma_{d}(\boldsymbol{\theta})=\frac{1}{n}d_{i}(\boldsymbol{\theta})d_{i}(\boldsymbol{\theta})^{\top}$. ∎
A.3 Proof of Lemma 1
Proof.
Case I. Baseline model $f_{0}$ is misspecified, propensity score model $e$ is correctly specified.
$\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})]$ =(A1)${{=}}\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{int}_{j},\mathbf{X}^{out}_{i},Y_{i}(\bar{a}_{i})]]$
Consider a particular summand of $\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})]$,
$\mathbb{E}[\mathbb{E}[\phi_{i}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})$ $\mid\mathbf{X}^{int}_{j},\mathbf{X}^{out}_{i},Y_{i}(\bar{a}_{i})]]$
$=\mathbb{E}[\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})(Y_{i}-\mu(\bar{a}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\theta}_{0}))$
$\quad\quad\quad\quad\quad\cdot\mathbb{E}[\sum_{j=1}^{J}H_{ij}\mathbb{E}[(A_{j}-\pi(\mathbf{X}^{int}_{j},\boldsymbol{\gamma}_{0}))\mid\mathbf{X}^{int}_{j},\mathbf{X}^{out}_{i},Y_{i}(\bar{A}_{i}),\mathbf{H}_{i}]]]$
Since $\pi$ is correctly specified, the inner expectation evaluates to 0.
Case II. Baseline model $f_{0}$ is correctly specified, propensity score model $p$ is incorrectly specified.
$\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})]$ $=\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J}]]$
=(A3)${{=}}\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\bar{A}_{i}]]$
=(A4)${{=}}\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]]$
Consider a particular summand:
$\mathbb{E}[\mathbb{E}[\phi_{ni}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})$ $\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]]$
$=\mathbb{E}[\lambda_{i}(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})(\bar{A}_{i}-\widehat{\bar{A}}_{i})$
$\quad\quad\quad\quad\quad\cdot\mathbb{E}[\mathbb{E}[(Y_{i}-\mu(\bar{a}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\theta}_{0}))\mid\mathbf{X}^{out}_{i},\bar{A}_{i},\mathbf{H}_{i}]]]$
Since $\mu$ is correctly specified, the inner expectation evaluates to 0. ∎
A.3.1 Proof for Theorem 1
To perform inference of $\mathbb{E}[Y_{i}(A;\boldsymbol{\theta}_{0})]$ under A Learning, we will require the following assumptions:
Inference Assumptions
  1. (A11)
    As $n,J\rightarrow\infty$, $\frac{J}{n}\rightarrow\mathcal{R}$ a fixed constant.
  2. (A12)
    $\Sigma_{\beta,\alpha}$, $\Sigma_{\phi}$ and $\Sigma_{\gamma}$, defined below in Section A.3.1, is finite and has a non degenerate limit
  3. (A13)
    $\Omega_{\psi}$, defined below in Section A.3.1, is positive semi-definite.
  4. (A14)
    Regularity of parameter space. Parameters $\boldsymbol{\gamma}$ lie in a compact space $\boldsymbol{\gamma}$ with $\boldsymbol{\gamma}_{0}$ as an inner point.
  5. (A15)
    The score function for $A_{j}$, $\phi_{\gamma}$, is continuously differentiable in $\mathbf{X}^{int}_{j}$ and $\boldsymbol{\gamma}$, and dominated by an integrable function.
  6. (A16)
    $V_{d}=\lim\limits_{J\rightarrow\infty}\frac{1}{J}\sum_{j=1}^{J}[-\frac{\partial\phi_{\boldsymbol{\gamma}}(A_{j};\boldsymbol{\gamma})}{\partial\boldsymbol{\gamma}}|_{\boldsymbol{\gamma}=\boldsymbol{\gamma}_{0}}]$ is finite and non degenerate
  7. (A17)
    $V_{\phi}=\lim\limits_{J\rightarrow\infty}\frac{1}{J}\sum_{j=1}^{J}\phi(A_{j};\boldsymbol{\gamma}_{0})\phi(A_{j};\boldsymbol{\gamma}_{0})^{\top}$ is positive semi-definite.
Assumptions (A12)-(A13) are for a well-defined covariance matrix similar to (A9)-(A10) in Section 2.3.1. Assumptions (A14)-(A17) are primarily for showing Lemma 2, which shows we have asymptotic normality of estimating $\boldsymbol{\gamma}_{0}$. We still require (A7)-(A8) for consistency of our estimates.
Proof.
Let $\Phi_{n}^{\mathrm{aug}}$ be $\Phi_{n}$ augmented with Equation 11. For example, if $f_{0}$ is linear in $\mathbf{X}^{out}_{i}$, we will augment with $(1,\mathbf{X}^{out}_{i})$. I.e.
$\Phi_{n}^{\mathrm{aug}}$ $=\begin{bmatrix}\sum_{i=1}^{n}&\frac{\partial}{\partial\boldsymbol{\alpha}}f_{0}(\mathbf{X}^{out}_{i},\boldsymbol{\alpha})(Y_{i}-\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\alpha},\boldsymbol{\beta}))\\ \sum_{i=1}^{n}\sum_{j=1}^{J}&\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})(Y_{i}-\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\alpha},\boldsymbol{\beta})))\\ &\quad\quad\quad\quad\quad\cdot\frac{1}{J}H_{ij}(A_{j}-\pi(\mathbf{X}^{int}_{j},\boldsymbol{\gamma}))\end{bmatrix}$ (A.1)
Take a Taylor expansion to find
$\Phi_{n}^{\mathrm{aug}}(\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\gamma}})$ $=\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})+\frac{\partial\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta},\boldsymbol{\alpha},\boldsymbol{\gamma}_{0})}{\partial(\boldsymbol{\beta}^{\top},\boldsymbol{\alpha}^{\top})}|_{\boldsymbol{\beta}=\boldsymbol{\beta}_{0},\boldsymbol{\alpha}=\boldsymbol{\alpha}_{0}}\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix}$
$\quad\quad\quad+\frac{\partial\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma})}{\partial\boldsymbol{\gamma}^{\top}}|_{\boldsymbol{\gamma}=\boldsymbol{\gamma}_{0}}(\hat{\boldsymbol{\gamma}}-\boldsymbol{\gamma}_{0})+\frac{1}{2}T_{2}+R_{nJ}$
where
$T_{2}$ $=\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix}^{\top}\frac{\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta},\boldsymbol{\alpha},\boldsymbol{\gamma}_{0})}{\partial(\boldsymbol{\beta}^{\top},\boldsymbol{\alpha}^{\top})\partial(\boldsymbol{\beta},\boldsymbol{\alpha})}|_{\boldsymbol{\beta}=\boldsymbol{\beta}_{0},\boldsymbol{\alpha}=\boldsymbol{\alpha}_{0}}\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix}$
$\quad\quad\quad+2\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix}^{\top}\frac{\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta},\boldsymbol{\alpha},\boldsymbol{\gamma})}{\partial(\boldsymbol{\beta}^{\top},\boldsymbol{\alpha}^{\top})\partial(\boldsymbol{\gamma}^{\top})}|_{\boldsymbol{\beta}=\boldsymbol{\beta}_{0},\boldsymbol{\alpha}=\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}=\boldsymbol{\gamma}_{0}}(\hat{\boldsymbol{\gamma}}-\boldsymbol{\gamma}_{0})$
$\quad\quad\quad+(\hat{\boldsymbol{\gamma}}-\boldsymbol{\gamma}_{0})^{\top}\frac{\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma})}{\partial\boldsymbol{\gamma}^{\top}\partial\boldsymbol{\gamma}}|_{\boldsymbol{\gamma}=\boldsymbol{\gamma}_{0}}(\hat{\boldsymbol{\gamma}}-\boldsymbol{\gamma}_{0})$
and
$R_{nJ}$ $=||\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix},\hat{\boldsymbol{\gamma}}-\boldsymbol{\gamma}_{0}||^{3}\frac{\sqrt{2}^{3}}{3!}\cdot\sup_{(\boldsymbol{\beta},\boldsymbol{\alpha}),\boldsymbol{\gamma}}|\frac{\partial^{3}}{\partial^{a}(\boldsymbol{\beta},\boldsymbol{\alpha})\partial^{b}\boldsymbol{\gamma}}\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta},\boldsymbol{\alpha},\boldsymbol{\gamma})|,\quad a+b=3$
Under regularity conditions (A7)-(A8) and (A14)-(A15), we have consistency of $\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\gamma}}$ and boundedness of these derivatives. Hence, $T_{2}=R_{nJ}=o_{P}(\sqrt{n})$, noting that remainder terms corresponding to $\boldsymbol{\gamma}$ are $o_{P}(\sqrt{J})=o_{P}(\sqrt{n})$ by (A11).
Rearranging, we have
$\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix}=$
$\quad\quad(\frac{\partial\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta},\boldsymbol{\alpha},\boldsymbol{\gamma}_{0})}{\partial(\boldsymbol{\beta}^{\top},\boldsymbol{\alpha}^{\top})}\Bigr{|}_{\begin{subarray}{c}\boldsymbol{\beta}=\boldsymbol{\beta}_{0}\\ \boldsymbol{\alpha}=\boldsymbol{\alpha}_{0}\end{subarray}})^{-1}[\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})+\frac{\partial\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma})}{\partial\boldsymbol{\gamma}^{\top}}\Bigr{|}_{\boldsymbol{\gamma}=\boldsymbol{\gamma}_{0}}(\hat{\boldsymbol{\gamma}}-\boldsymbol{\gamma}_{0})]$
$\quad\quad\quad\quad+o_{p}(1/\sqrt{n})$
Let $\Sigma_{\beta,\alpha}=\lim\limits_{n\rightarrow\infty}\frac{1}{n}\frac{\partial\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta},\boldsymbol{\alpha},\boldsymbol{\gamma}_{0})}{\partial(\boldsymbol{\beta}^{\top},\boldsymbol{\alpha}^{\top})}\Bigr{|}_{\begin{subarray}{c}\boldsymbol{\beta}=\boldsymbol{\beta}_{0}\\ \boldsymbol{\alpha}=\boldsymbol{\alpha}_{0}\end{subarray}}$ and $\Sigma_{\gamma}=\lim\limits_{n\rightarrow\infty}\frac{1}{n}\frac{\partial\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma})}{\partial\boldsymbol{\gamma}^{\top}}\Bigr{|}_{\boldsymbol{\gamma}=\boldsymbol{\gamma}_{0}}$.
Invoking (A12), we have
$\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix}$ $=\Sigma_{\beta,\alpha}^{-1}[\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})+\Sigma_{\gamma}(\hat{\boldsymbol{\gamma}}-\boldsymbol{\gamma}_{0})]+o_{p}(1/\sqrt{n})$
Then we have,
$\sqrt{n}\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix}=\Sigma_{\beta,\alpha}^{-1}[\frac{1}{\sqrt{n}}\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})+\Sigma_{\gamma}\sqrt{n}(\hat{\boldsymbol{\gamma}}-\boldsymbol{\gamma}_{0})]+o_{p}(1)$
Let a summand of $\Phi_{n}^{\mathrm{aug}}$ be denoted by
$\phi_{i}^{\mathrm{aug}}$ $=\begin{bmatrix}&\frac{\partial}{\partial\boldsymbol{\alpha}}f_{0}(A;\mathbf{X}^{out}_{i},\boldsymbol{\alpha})(Y_{i}-\mu(A;\mathbf{X}^{out}_{i},\mathbf{H}_{i},\boldsymbol{\alpha},\boldsymbol{\beta}))\\ &\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})(Y_{i}-\mu(A;\mathbf{X}^{out}_{i},\mathbf{H}_{i},\boldsymbol{\alpha},\boldsymbol{\beta})))\\ &\quad\quad\quad\quad\quad\cdot\frac{1}{J}\sum_{j=1}^{J}H_{ij}(A_{j}-\pi(\mathbf{X}^{int}_{j},\boldsymbol{\gamma}))\end{bmatrix}$ (A.2)
Then, by Lemma 2,
$\sqrt{n}\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix}$ $=\Sigma_{\beta,\alpha}^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi_{i}^{\mathrm{aug}}+\Sigma_{\beta,\alpha}^{-1}\Sigma_{\gamma}\frac{1}{\sqrt{\mathcal{R}}}\frac{1}{\sqrt{J}}\sum_{j=1}^{J}\varepsilon_{j}+o_{p}(1)$
Note that conditioning on $\mathbf{X}^{out}_{i},\mathbf{X}^{int}_{j},\mathbf{H}_{i},A$, each term on the RHS is mean 0.
Let $\Sigma_{\phi}=\lim\limits_{n\rightarrow\infty}\sum_{i=1}^{n}\phi_{i}^{\mathrm{aug}}(\phi_{i}^{\mathrm{aug}})^{\top}$, and define
$\Omega_{\phi}=(\Sigma_{\beta,\alpha}^{-1})\Sigma_{\phi}(\Sigma_{\beta,\alpha}^{-1})^{\top}$ (A.3)
$\Omega_{\gamma}=(\frac{1}{\sqrt{\mathcal{R}}}\Sigma_{\beta,\alpha}^{-1}\Sigma_{\gamma})\Omega_{\varepsilon}(\frac{1}{\sqrt{\mathcal{R}}}\Sigma_{\beta,\alpha}^{-1}\Sigma_{\gamma})^{\top}$ (A.4)
Then we can apply the multiplier central limit theorem 2 to each term to conclude asymptotic normality with variance covariance given by
$\Omega=\Omega_{\phi}+\Omega_{\gamma}$ (A.5)
In practice, we need to estimate $\mathcal{R}$ using $\frac{J}{n}$. By Slutsky’s, the result follows. ∎
A.4 Propensity Score Helper Lemma
For Theorem 1, we require consistent estimation and asymptotic normality of propensity score model $\pi$. We state this in the following helper lemma.
Lemma 2.
Assume (A7), (A11), (A14)-(A17). Then we have consistent estimation and asymptotic normality of propensity score model $\pi$. I.e.
[…]
where for $\Omega_{\varepsilon}=\lim\limits_{J\rightarrow\infty}\sum_{j=1}^{J}\varepsilon_{j}\varepsilon_{j}^{\top}$ and $\varepsilon_{j}$ independent mean zero random vectors.
Proof.
The proof follows identically to Section 2.3.1. ∎
A.5 Additional Helper Results
The following is Lemma 10.5 from (Kosorok, 2008), restated here for convenience.
Theorem 2 (Multiplier Central Limit Theorem).
[…]
Appendix B Descriptive Statistics
We compute basic descriptive statistics on our outcome and intervention dataset.
Variable Mean Range
Total NOx controls 2.94 (0, 24)
log(Heat input) 14.45 (8.98, 17.32)
log(Operating time) 7.20 (5.39, 8.93)
% Operating capacity 0.64 (0.07, 1.17)
% Selective non-catalytic reduction 0.26 [0, 1]
ARP Phase II 0.71 {0, 1}
Scrubbed 0.19 {0, 1}
Table B.1: Summary of intervention level covariates from power plant data.
Variable Mean Range
% White 0.89 (0, 1)
% Female 0.56 (0, 1)
% Urban 0.38 (0, 1)
% High school graduate 0.36 (0, 1)
% Poor 0.13 (0, 1)
% Moved in last 5 years 0.42 (0, 1)
% Households occupied 0.87 (0.015, 1)
% Smoke 0.26 (0.13, 0.44)
Mean Medicare age 74.94 (68.96, 96.26)
Mean temperature (K) 287.69 (275.34, 301.14)
Mean relative humidity (%) 0.0087 (0.0045, 0.016)
log(Population) 8.16 (1.39, 11.65)
Population per Square Mile 1123.78 (0.067, 158,503.38)
IHD Hosp. Rate (per 10k PY) 0.29 (0, 0.35)
IHD Hosp. Count 22.50 (0, 557)
Table B.2: Summary of outcome level covariates from Medicare Beneficiary data. The total number of IHD Hospitalizations in the dataset (2005) are 600,346 and the total rate of IHD Hospitalizations per 10k person-years is 775.362.
Appendix C Additional Simulation Details
C.1 Simulation Parameters
The simulation parameters are found in the supplement code, and summarized below:
$\boldsymbol{\theta}=($ $-0.000955,0.0288,0.0382,-0.000148,-0.00227,0.0167,-0.0199,$
$0.0396,0.0152,0.0173,-0.0119,0.0161,0.0329,0.0365,$
$-0.0203,0.0221,-0.0181,-0.0262,2.170e-05,0.0305,-0.0310,$
$0.0357,-0.0187,0.00968,0.0204,0.0269,0.00420,-0.000470,$
$-0.000344,-0.000490,0.000127,0.00115,0.00140,0.00118,0.00133,$
$0.00120,-0.000423,-0.000867,0.000361,-0.00135,-0.001362,5.567e-05,$
$0.000982,-0.000606,0.000586,-0.00121,-0.000864,-0.000517,$
0.00135,0.000558,0.00103,0.00106,0.00117,0.000676-0.00135,-0.000558,0.00103,0.00106,0.00117,-0.000676)- 0.00135 , - 0.000558 , 0.00103 , 0.00106 , 0.00117 , - 0.000676 )
$\boldsymbol{\gamma}=($ $-0.681,0.131,-0.704,0.386,0.334,0.424,$
0.00141,0.00471,0.010,0.0140,0.0107,1.449 0.00141,-0.00471,0.010,-0.0140,-0.0107,-1.449)0.00141 , - 0.00471 , 0.010 , - 0.0140 , - 0.0107 , - 1.449 )
C.2 Simulation Metrics
The specific formulas for the metrics in the simulation are as follows:
The parameter error for the treatment effect function $f_{A}$ is given by
$\texttt{Bias}=||\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}_{0}||_{2}$ (C.1)
and the root mean squared error in the estimated treatment effect of interest is given by
$\texttt{RMSE}=\mbox{RMSE}(\widehat{\texttt{TotalEffect}},\texttt{TotalEffect})$ (C.2)
Appendix D Data Preprocessing
D.1 Intervention Unit Trimming
We trim intervention units with low treatment probabilities. The histogram for fitted intervention unit probabilities are shown in Section D.1.
D.2 Cost Model Training
We select our model based on Normalized Mean Absolute Error (NMAE). Table D.1 shows the NMAE for each cost model. For a model $\mathcal{M}$ with predictions $\hat{C}_{j}=\mathbb{E}[C_{j}\mid\mathbf{X}^{int}_{j};\mathcal{M}]$, NMAE is calculated as follows:
$NMAE(\mathcal{M})=\frac{1}{J}\sum_{j=1}^{J}\frac{|C_{j}-\hat{C}_{j}|}{|\hat{C}_{j}|}$ (D.1)
Cost Model Results
$\mathcal{M}$ NMAE
LM $0.91$
SVM $0.55$
RF $0.54$
Table D.1: Cost model NMAE on the validation set.
The random forest model performs best on our validation dataset, so random forest is used to generate cost predictions. The variable importance plot (under Increase in Node Purity) is shown in Section D.2. We find Total SO2 Emissions and and Total Load yield high variable importance for predicting cost, which makes sense given the scientific domain.
In Fig. D.1, we display the random forest model predictions against the true values, on the training and validation set.
Refer to caption Refer to caption
(a) (b)
Figure D.1: (a) displays the predictions against truth in our training set and (b) displays the predictions against the truth in our validation set
Appendix E Additional Analysis Details
E.1 Exploratory Analysis of $\widehat{\texttt{TotalEffect}}$
We perform an exploratory analysis of $\texttt{TotalEffect}_{j}$ testing $H_{0}:\texttt{TotalEffect}_{j}\geq 0,H_{a}:\texttt{TotalEffect}_{j}<0$. The p-values for each of the $J$ power plants are shown in Fig. E.1. Approximately 95% of p-values are $<0.05$. Caution should be taken in interpreting these findings, as dependence and false discovery should be accounted for in a formal statistical test. Regardless, these findings support that there are statistically significant population treatment effects with each scrubber installation.
Figure E.1: P-Values from testing $H_{0}:\texttt{TotalEffect}_{j}\geq 0,H_{a}:\texttt{TotalEffect}_{j}<0$.
E.2 Nonlinear Policy Analysis
In Section 4, we performed policy analysis using linear models akin to Equation 15, Equation 16. In this section, we present a brief overview of results using nonlinear models. As shown in Table E.1, the potential reduction in hospitalization rates and counts increases notably with more complex models. As stated in the main text, we report the linear model results for interpretability and to err on the conservative side.
Estimated Reductions
Model Rates Counts
Linear 44.51 31,906
Quadratic 58.50 58,984
Cubic 66.58 76,761
Sin/Cos 67.28 67,388
Table E.1: Estimated reductions in IHD hospitalization rates per 10k person-years and IHD hospitalization counts, varying the model link, with no budget constraints.
E.3 Count Analysis
The following are figures akin to Fig. 2, but for IHD Hospitalization Counts. The counts are obtained by multiplying the rates by the number of person-years observed for each zip code. As shown in Fig. E.2, $>99\%$ of power plants are protective. Many of the plants that are most protective in terms of rates are also protective in terms of counts. However, there are more plants that are similarly protective in the northeast and southeastern US. Similar to the cost-constrained rates analysis, much of the gain is experienced from 10% – 30% of universal scrubber treatment cost. In addition, from 60%, much of the gain stabilizes.
Refer to caption Refer to caption
(a) (b)
Figure E.2: (a) displays the plot of $\widehat{\texttt{TotalEffect}}_{j}$ on a US map, colored by intensity of $\widehat{\texttt{TotalEffect}}_{j}$. (b) displays the histogram of $\widehat{\texttt{TotalEffect}}_{j}$.
Refer to caption Refer to caption Refer to caption
(a) 10% Total Cost (b) 20% Total Cost (c) 30% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 40% Total Cost (b) 50% Total Cost (c) 60% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 70% Total Cost (b) 80% Total Cost (c) 90% Total Cost
Figure E.3: Grid with the reduction of IHD Hospitalizations, varying the spending from 10%-90% of budget
Appendix F Model Generalizations and Extensions
Multiple Timepoints
If we have $n$ samples over $t\in[T]$ timepoints, we may follow the same approach as in Schulte et al. (2014). For illustration purposes, we once again assume the outcome model at each timepoint factorizes by intervention unit as in Equation 5:
$Y_{it}(\bar{a}_{it})$ $=\mu_{t}(\bar{a}_{it};\mathbf{X}_{it}^{out},\mathbf{H}_{i},\boldsymbol{\theta}_{0t})+\epsilon_{it}$
$=f_{0t}(\mathbf{X}_{it}^{out},\boldsymbol{\alpha}_{0t})+\bar{a}_{it}f_{At}(\mathbf{X}_{it}^{out},\boldsymbol{\beta}_{0t})+\epsilon_{it}$ (F.1)
Above, each quantity is the analog as described in Equation 5, but indexed a second time for timepoint $t\in[T]$.
Q-Learning
We propose to solve the following estimating equation backwards in time $t=T,T-1\dots 1$.
$\frac{1}{n}\sum_{i=1}^{n}\phi_{t}(Y_{it};\boldsymbol{\theta})=0$ (F.2)
where $\phi_{t}(Y_{it};\boldsymbol{\theta}_{t})=(Y_{it}-\mu_{t}(\bar{A}_{it};\mathbf{X}_{it}^{out},\mathbf{H}_{i},\boldsymbol{\theta}_{t}))d_{it}(\boldsymbol{\theta}_{t})$, $d_{it}(\boldsymbol{\theta}_{t})=\frac{\partial\mu_{t}(\bar{A}_{it},\mathbf{X}_{it},\mathbf{H}_{i},\boldsymbol{\theta}_{t})}{\partial\boldsymbol{\theta}_{t}}$.
Specifically, at any timepoint $t$, we obtain $\hat{\pi}^{*}_{t}(\widehat{\texttt{TotalEffect}_{jt}})$. Thus at time $t-1$, we have the optimal policy value assuming policymakers make the optimal policy choice at time $t$. With backwards induction, model-correctness, and straightforward extensions of the assumptions mentioned throughout Section 2, this strategy can be shown to be valid.
A-Learning
The result follows similarly to above. Extending the estimating equations in question Equation 11, Equation 12, Equation 10 to $T$ timepoints and solving for these backwards in time yields a multi-time point extension to A Learning.
Nonparametric Models
The approaches above can be extended to more flexible models assuming nuisance parameters are estimated at $\sqrt{n}$ consistent rate. Valid inference would follow from Section 2.3.1 and Theorem 1 since higher order terms would decay to 0 fast enough. In particular, one may employ nonparametric series estimators for functions that satisfy relevant smoothness conditions (Giné & Nickl, 2015).
Mean Outcome Form
The proposed framework permits use of more general mean outcome models than that considered in Equation 5. Equation 5 enables us to obtain closed-form or efficient solutions to Equation 4. In general, we may require higher computational burden depending on the assumed outcome form. Interpretability also goes down since we would not be able to directly interpret TotalEffect in terms of the outcome model as in Equation 6.